Find the solution in point-slope form of the the line perpendicular to 3x+2y=3y-7 and passing through (3,-2)
Answer by lwsshak3(11628)

Find the solution in point-slope form of the the line perpendicular to 3x+2y=3y-7 and passing through (3,-2)
***
3x+2y=3y-7
y=3x+7
slope,m=3
slope of perpendicular line=-1/3(negative reciprocal)
y-(-2)=(-1/3)(x-3)
y+2=-(x-3)/3
3y+6=-x+3
x+3y=-3